6. Consider the intial value problem: 494 + 42y + 9y = 0, y(0) = a > 0, y(0) = -1. a. Find the solution in terms of a. Giv
Answer #1

49 yl + 42 y + 9y = 0 ( 9 ) Aunilory equation Ylo) = aro & yro)=-1 49 m2 +42m +9=0 m= -42 (42)² - 36 (49) 2x 49 -42 1 0 98 -